A right triangle has a perimeter of 26 inches.The length of each side is increased to 7 times it original length. What is the perimeter of the larger triangle in inches?
Answer by jim_thompson5910(35256)

Call the sides: s1, s2, s3


The perimeter is P = s1 + s2 + s3


The given perimeter is P = 26, so 26 = s1 + s2 + s3


Each side s1, s2, s3 is "increased to 7 times it original length", so we have the new sides 7*s1, 7*s2, 7*s3.


They add up to: 7*s1 + 7*s2 + 7*s3 = 7*(s1 + s2 + s3) = 7*P = 7*26 = 182


So the new perimeter of the larger triangle is 182 inches